The midpoint of 0 up to, not including 6 class is

  1. 2
  2. 4
  3. 2.5
  4. 0
Answer #1

The midpoint of 0 up to, not including 6 class

That is, 0 up to 5.

Let,

midpoint = (upper class - lower class)/2

Midpoint = (5 - 0)/2 = 2.5

Midpoint = 2.5

Answer : c) 2.5
Answer #2

The midpoint of the class interval from 0 up to, but not including 6 can be calculated by taking the average of the endpoints.

The endpoints are 0 and 6. To find the average, we add the two values and divide by 2:

(0 + 6) / 2 = 6 / 2 = 3

Therefore, the midpoint of the class interval from 0 up to, but not including 6 is 3.

The correct answer is: c) 2.5


answered by: mervetokaz